OpenStudy (anonymous):

Find two Square roots for 4i. Leave in standard/rectangular form. Help would be much appreciated! :)

OpenStudy (anonymous):

\[\sqrt4 = \pm2\]

OpenStudy (anonymous):

thanks, but that's for 4, not 4i :)

OpenStudy (anonymous):

\[4i= 4 e ^{i \Pi/2}\]

OpenStudy (anonymous):

so, square root of 4i is \[\pm2 e ^{i \Pi/4}=\pm2(\cos \Pi/4 + i \sin \Pi/4)\]

OpenStudy (anonymous):

simplify it, u will get: \[\pm \sqrt{2} (1+i)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome :)